OpenStudy (anonymous):

express the following questions in slope intercept form:5x+8y=40

OpenStudy (kc_kennylau):

That's the same as putting y as the subject :)

OpenStudy (kc_kennylau):

5x+8y=40 8y=40-5x y=-5/8x+40/8 y=-0.625x+5

OpenStudy (anonymous):

slope intercept form is y=mx+b

OpenStudy (kc_kennylau):

m=-0.625 b=5

OpenStudy (anonymous):

Lexi answered 5 months ago y=mx+b y is the answer so it stays y. m is the slope x is the x intercept ehich stays x b is the y intercept so the form you see it in now is ax+by=c so to solve for y subtract ax on both sides then divide by whatever to make by equal to one sometimes no need to divide. here is an example 6x+5y=-30 subtrract 6x from both sides 5y=-6x-30 divide by negative five y=6/5x+6 hope that helped i love algebra!!:)